Project

General

Profile

20170411 » History » Version 1

Boyd, Steven, 09 April 2017 16:47

1 1 Boyd, Steven
h1. Ops/Comm Scheduling meeting : 11/4/2017
2
3
Location : CR21 in R1 at 14:00 pm
4
5
Phone access :  *New* phonebridge number
6
UK: 08082380274
7
US: 631.621.5253
8
Access code: 8161978083
9
Leader code: 2735
10
11
h2. Agenda
12
13
* Actions
14
** Ajit - No further development of Hanlet SMS system. Need to propose a basic messaging system still to alleviate need for 24Hr monitoring in person. - on hold
15
*** N.B. that we already have an alarm display in the ISIS MCR. CWhyte to check with ABross/JBoehm that automated texting would be acceptable for removing requirement for continuous oversight while CC magnets powered.
16
*** We have the SMS to email gateway, including monitoring for e-mail responses, but a lot of work remains in driving this from the Alarm Handler.
17
** Macwaters & Boehm - Arrange to install new Sumitomo compressor adsorbers? - stands
18
** Macwaters - Re-make dodgy compressor2 remote cable connector - stands
19
** Macwaters - Prepare new F70 compressors for use. Remove old overheated CSW71 to R9. Speak to Sumitomo on April 5th regards how to repair. - in progress
20
*** New F70s on site - RHale to fit mains leads, then move to Hall?
21
** Macwaters - Replace old vac sensors and new valve on cryo 1&2 - in progress
22
** Ajit - Decide on new way to implement what was MOM/DC state machine override function - in progress
23
*** Done for DS, needs re-coding for the rest; also need documentation and training for operators.
24
** Macwaters - Report on status of FC psu repair (stands)
25
*** Has arrived at Sorenson and been examined, but no feedback to us as yet 
26
** Hall Manager - water system change-over procedure (stands)
27
** EOverton - Status of @miceraid5@ disk cleanup (stands)
28
*** Was done for 2016/05, needs checking for 2017/01
29
** MUchida - investigate moving Tracker waveguides around to cure "hole". Is Ed happy for this to progress? Yes. (stands)
30
** MCourthold - show JBoehm how to use the Picologger for quench data. Craig can help. (stands)
31
** JBoehm & Macwaters - get the West Wall compressors correctly labelled, as various have been re-purposed (stands)
32
33
** CWhyte to check with ABross/JBoehm that automated texting would be acceptable for removing requirement for continuous oversight while CC magnets powered.
34
** CWhyte to contact DL to clarify their [[Hall_Task_List_Spring17|Hall Task List]] items
35
** PFranchini to go through servers and turn off IPMI in BIOS
36
37
* DC update
38
* LH2 update
39
* [[Hall_Task_List_Spring17|List of Hall activities in the shutdown]] (Colin) 
40
41
AOB:
42
43
* Last meeting : [[20170328|Tuesday Apr 4 2017]]
44
* Next meeting : [[20170411|Tuesday Apr 25 2017]]